# Quantifying Very Extreme Precipitation and Temperature Using Huge Ensembles Generated by Machine Learning-based Climate Model Emulators

### v1

@@ 3.2 Temperature
Figure 7: Maximum daily temperature from emulator for different numbers of simulated years compared to maximum in ERA5 (1940-2022). One-to-one lines are in red.
Fig. 8 provides information about the extremes climatology (estimated AEP temperatures and shape parameters) in the ensemble, based on the threshold exceedances approach using $n=499$ exceedances. The AEP temperatures show spatial patterns expected based on U.S. temperature climatology with the most extreme temperatures seen in the more arid inland areas of the western U.S. As expected, the shape parameter estimates are almost all negative. Note that the local maximum seen in the north-central U.S. in Fig. 8 is likely from sampling variability (given the lack of topography in that area of the U.S.) and associated with the coinciding anomalously large shape parameter estimates seen in Fig. 8e.
Figure 8: Emulator-based extreme temperature climatology: (a) 1-in-1000 year AEP temperature estimates, (b) 1-in-100000 year AEP temperature estimates, (c) shape parameter estimates, with corresponding standard error estimates (d), (e), and (f), respectively.
Next we assess the performance of the GEV and POT extreme value fitting approaches for estimation for AEP temperatures corresponding to shorter time durations (1000 and 10000 years) to assess the performance of EVA. Fig. 9 shows that both GEV and POT (regardless of the threshold) produce AEP temperature estimates that are very similar to the empirical quantiles. This contrasts with the bias seen for precipitation; with the more bounded temperature distributions, it is not necessary to use longer blocks or larger thresholds to minimize bias. These results provide support for using the EVA-based AEP temperature estimates for longer durations (such as 100000 and one million years). We did notice that the numerical optimization for the GEV approach was more fragile than for POT, requiring more careful initialization to find the maximum of the log-likelihood at many locations.
@@
Figure 10: Comparison of 1-in-100000 year AEP temperature estimates based on analysis of full-year data to maximum of seasonally-stratified estimates. Results are shown for the second approach discussed in the methods. One-to-one lines are in red.
@@ 4 Discussion
A model- (i.e., simulation-) based approach carries numerous benefits. Central to these is the large set of samples that characterize variability, with each sample capturing realistic spatio-temporal structure. One can calculate any functional of interest (such as total precipitation over a particular watershed for a particular length of time) for each sample. The resulting samples of the functional characterize the variability of that quantity of interest, with no need for complicated statistical techniques to account for joint uncertainty and the effects of nonlinearity.
The major limitation is that the model/emulator must be fit for purpose, producing output similar enough to the real world for the estimates to be useful. This requires very careful assessment, assessment that is not part of this work and which for the most extreme events is by definition difficult or impossible. An additional limitation is the need for higher-resolution output in space and time, both for understanding impacts of interest and potentially for validating results from lower-resolution output, with sub-daily extremes being particularly challenging (NASEM, 2024). Finally, we note that the forcing conditions (from the years 2001-2022), of which the SSTs may be most critical, may not be sufficiently representative and may constrain the simulated weather produced.

### v2

@@ 3.2 Temperature
Figure 7: Maximum daily temperature for each grid cell over the specified time period from emulator for different numbers of simulated years compared to maximum in ERA5 (1940-2022). One-to-one lines are in red.
Fig. 8 provides information about the extremes climatology (estimated AEP temperatures and shape parameters) in the ensemble, based on the threshold exceedances approach using $n=499$ exceedances. The AEP temperatures show spatial patterns expected based on U.S. temperature climatology with the most extreme temperatures seen in the more arid inland areas of the western U.S. As expected, the shape parameter estimates are almost all negative.
Figure 8: Emulator-based extreme temperature climatology: (a) 1-in-1000 year AEP temperature (deg. C) estimates, (b) 1-in-100000 year AEP temperature (deg. C) estimates, (c) shape parameter estimates, with corresponding standard error estimates (d), (e), and (f), respectively. The local maximum seen in the north-central U.S. (southern Minnesota), an area with little topographic variation, appears to be from sampling variability and not optimization problems and is associated with coinciding anomalously large shape parameter estimates.
Next we assess the performance of the GEV and POT extreme value fitting approaches for estimation for AEP temperatures corresponding to shorter time periods (1000 and 10000 years) to assess the performance of EVA. Fig. 9 shows that both GEV and POT (regardless of the threshold) produce AEP temperature estimates that are very similar to the empirical quantiles. This contrasts with the bias seen for precipitation; with the more bounded temperature distributions, it is not necessary to use longer blocks or larger thresholds to minimize bias. These results provide support for using the EVA-based AEP temperature estimates for longer periods (such as 100000 and one million years). We did notice that the numerical optimization for the GEV approach was more fragile than for POT, requiring more careful initialization to find the maximum of the log-likelihood at many locations.
@@
Figure 10: Comparison of 1-in-100000 year AEP temperature estimates based on analysis of full-year data to aggregation over seasonally-stratified estimates. Results are shown for the second approach discussed in the methods. One-to-one lines are in red.
@@ 4 Discussion
A model- (i.e., simulation-) based approach carries numerous benefits. A central one is the large set of samples that characterize variability, with each sample capturing realistic spatio-temporal structure. One can calculate any function of interest (such as total precipitation over a particular watershed and time duration) for each sample. The resulting samples of the function characterize the variability of that quantity, with no need for complicated statistical techniques to account for joint uncertainty and the effects of nonlinearity.
The major limitation is that the model/emulator must be fit for purpose, producing output similar enough to the real world for the estimates to be useful. This requires very careful assessment, assessment that is not part of this work and which for the most extreme events is by definition difficult or impossible. An additional limitation is the need for higher-resolution output in space and time, both for understanding impacts of interest and potentially for validating results from lower-resolution output, with sub-daily extremes being particularly challenging (NASEM, 2024). Finally, we note that the forcing conditions (from the years 2001-2022), particularly the SSTs, may not be sufficiently representative and may constrain the simulated weather produced.
@@
Much larger sample sizes than 10560 will likely be available in the near future, raising the possibility of simply using empirical quantiles rather than EVA. However, empirical quantiles are also statistical estimates with associated uncertainty, and improvements in emulator efficiency will likely be leveraged at least in part towards emulation at higher resolution or for multiple scenarios, suggesting a need for EVA for the foreseeable future.